# Trust in Autonomous Human–Robot Collaboration: Effects of Responsive Interaction Policies

### v1

@@ 2.6 Analytic Strategy
Because the study relied fundamentally on spoken-language collaboration, analyses were structured to explicitly account for interaction viability. Sessions were classified as non-viable when severe communication breakdown prevented sustained dialogue between the participant and the robot, rendering the experimental manipulation inoperative.
Communication viability was operationalized using a dialogue-level metric derived from manual coding of system logs. For each session, the proportion of dialogue turns affected by speech-recognition failure (i.e., fragmented or unintelligible utterances) was computed. Sessions in which more than 60% of dialogue turns were affected were classified as non-viable, reflecting cases in which spoken-language interaction could not be meaningfully sustained. This criterion closely aligned with sessions independently flagged by the researcher during administration. It should be noted that because communication viability may itself be influenced by interaction policy, excluding non-viable sessions may have introduced selection bias and inflated estimated policy effects.
For this reason, analyses were conducted using three complementary approaches. Primary analyses were performed on an eligible sample excluding non-viable sessions, reflecting interactions in which the spoken-language protocol and experimental manipulation operated as intended. Full-sample analyses including all sessions were conducted as sensitivity checks. Finally, mechanism-focused analyses compared viable and non-viable sessions on interaction-process metrics (e.g., ASR failure rates, dialogue turn completion, task abandonment) to characterize how severe communication breakdown alters interaction dynamics. Trust measures from non-viable sessions were not interpreted as valid estimates of human–robot trust under functional interaction, as the robot was unable to sustain dialogue or collaborative behavior in these cases.
All analyses were conducted using R (version 4.5.1) within the Quarto framework. Data manipulation and visualization utilized the tidyverse suite of packages [40], with mixed-effects models fitted using the lme4 and lmerTest packages [3, 23] while Bayesian hierarchical models were fitted using the brms package [6]. Summary tables were generated using the gtsummary package [36]. All code used for data processing and analysis is available at: GitHub Repository
@@ 3 Results
Of the 29 completed sessions, 5 met the pre-specified criterion for non-viable interaction due to severe and persistent communication failure. These sessions were characterized by high rates of ASR failure, incomplete dialogue sequences, and skipped task stages.
Primary results are therefore reported for the eligible sample (n = 24), with full-sample and mechanism-focused analyses reported as sensitivity and exploratory analyses, respectively.
3.1 Primary Analysis: Eligible Sample
Simple descriptive comparisons of post-interaction trust measures indicated higher trust ratings in the responsive condition relative to the control condition across both trust scales (see Table 2). As indicated in Figure 4, average raw post-interaction scores on the Trust in Industrial Human–Robot Collaboration scale (TI-HRC) differed by approximately 26 points (Likert 1-5 converted to 0-100 scale for easier comparison across scales), while differences on the Trust Perception Scale–HRI (TPS-HRI) were approximately 15 points higher in the responsive condition compared to the control (Figure 4).
Importantly, overall task accuracy did not differ significantly between conditions (detailed task performance results are reported below), suggesting that observed differences in trust were not driven by differential task success but rather by the quality of the interaction process itself.
Table 2: Post-Interaction Trust Outcomes and Task Accuracy by Policy in the Eligible Sample
@@ 3.1 Primary Analysis: Eligible Sample
Analysis of dialogue interaction patterns confirmed successful manipulation of the interaction policies (see Table 3). Manual coding of dialogue transcripts revealed substantial and significant differences in robot behaviour across conditions (see Appendix C for coding scheme). In the responsive condition, the robot employed encouragement in 36% of dialogue turns compared to 0% in the control condition (p < .001), expressed empathy or acknowledged participant affect in 13% of turns versus 0% in the control (p < .001), and used collaborative language (e.g., “we,” “let’s”) in 42% of turns compared to 5% in control (p < .001). Interestingly, while both conditions included proactive check-ins following periods of participant silence, the control robot engaged in such check-ins at a higher rate (21% vs. 13% of turns, p = .033), consistent with its reactive policy that limited proactive assistance to structured silence monitoring.
Critically, proportions of communication breakdown did not differ between conditions (25% vs. 22%, p = .70), indicating that interaction policy did not systematically impact technical speech recognition viability among eligible participants. Participants in the responsive condition also exhibited higher levels of AI-detected engagement during interaction, with an average of 3.50 engaged responses (SD = 1.95) compared to 2.00 (SD = 2.21) in the control condition. Consistent with these dialogue differences, interactions in the responsive condition were characterized by longer session durations and slower robot response times, reflecting the latency and time associated with executing additional dialogue and affective support behaviours. As such, responsiveness in the present study reflects a bundled interaction profile rather than a single isolated manipulation. Observed trust differences may therefore partially reflect increased interaction exposure or dialogue richness in addition to affect-adaptive behavior per se. Future work should more finely disentangle these components.
@@ 3.2 Hierarchical Models
Bayesian models used weakly informative, scale-appropriate priors. The intercept was assigned a Normal(50, 25) prior, reflecting the 0–100 trust score range and centering estimates near the midpoint of the scale. Fixed effects were assigned Normal(0, 10) priors to constrain effects to plausible magnitudes. Random-effect and residual standard deviations were assigned Exponential(1) priors to regularize variance estimates while retaining flexibility. All models demonstrated satisfactory convergence ($\hat{R} \leq$ 1.01; effective sample sizes > 1000).
@@ 3.2.1 Primary Analysis: Eligible Sample (n = 24)
Bayesian estimation converged on similar effect magnitudes with high posterior certainty. As illustrated in Figure 5, for TI-HRC, the responsive condition showed a posterior median effect of $\beta$ = 14.98 (95% credible interval [7.29, 22.22]), with near-unity probability of a positive effect. For TPS-HRI, the posterior median was $\beta$ = 12.76 (95% credible interval [2.96, 22.06]), with posterior probability exceeding 99% that the effect was positive. Baseline NARS showed credible negative associations with both outcomes, while native English fluency showed negative associations that were credible for TPS-HRI but uncertain for TI-HRC. Model fit was substantial for TPS-HRI (conditional $R^{2}$ = 0.64) and moderate for TI-HRC (conditional $R^{2}$ = 0.42), with fixed effects explaining 16% and 21% of variance, respectively. The smaller item-level variance for TI-HRC suggests greater coherence among affective trust items under functional interaction conditions.
Figure 5: Posterior distributions of fixed effects from the final Bayesian mixed-effects models (eligible sample: n=24) predicting post-interaction trust. Fixed effects include Responsive vs Control policy, Negative Attitudes Towards Robots (NARS) and Non-native English speaker versus English native speaker. Half-eye densities show posterior distributions; points indicate posterior medians; thick and thin intervals denote 80% and 95% credible intervals, respectively. The dashed vertical line marks a null effect, and positive values indicate higher trust scores on the respective scale (points on a 0–100 scale).
@@ 3.2.3 Mechanism Analysis: Communication Breakdown (n=29)
To examine whether communication quality moderated the effect of interaction policy, Bayesian models were fitted in the full sample with the proportion of turns characterized by communication breakdown included as an interaction term. For TPS-HRI, the interaction between policy and communication breakdown was weak and centered near zero (posterior median $\beta$ = -1.14, 95% credible interval [-18.87, 16.28]), with only 55% probability of being negative and 50% probability of being significant. This broad, unstable posterior distribution suggests that evaluative trust judgments were relatively insensitive to graded variation in communication quality once a basic threshold of viability was established. In contrast, TI-HRC showed a more consistent pattern (see Figure 6). The interaction term showed a negative tendency (posterior median $\beta$ = -5.97, 95% credible interval [-23.01, 10.92]), with 76% probability of being negative, 71% probability of being significant, and 45% probability of being large. While responsive behaviour was associated with higher experienced trust under low levels of communication breakdown, this advantage diminished as communication failures accumulated. This differential sensitivity suggests that experienced trust depends critically on the robot’s sustained ability to engage in responsive interaction, whereas evaluative trust judgments may rely more on discrete moments of successful collaboration that can occur even within partially degraded interactions.
Figure 6: Predicted trust scores as a function of communication breakdown, stratified by interaction policy. Panels show experienced trust (left) and perceived trust (right). Lines represent posterior medians and shaded bands indicate 80% credible intervals. Communication breakdown is defined as the proportion of interaction turns flagged as breakdowns; predictions are shown at mean baseline negative attitudes toward robots.
@@ 3.4 Individual differences and correlational patterns
Consistent with the experimental manipulation, specific robot dialogue behaviours showed substantial correlations with trust outcomes. Robot empathy expression—quantified as the proportion of dialogue turns in which the robot acknowledged participant affect or expressed understanding—was strongly correlated with both experienced trust (TI-HRC; r = .53, p = .008) and perceived trust (TPS-HRI: r = .65, p = .001). Similarly, robot use of collaborative language (e.g., “we,” “let’s”) was positively associated with experienced trust (TI-HRC: r = .50, p = .012), as was robot encouragement (TI-HRC: r = .45, p = .027; TPS-HRI: r = .42, p = .040). These associations provide evidence that the specific affective and collaborative behaviours implemented in the responsive condition were linked to participants’ trust evaluations beyond simple condition assignment.
@@
Notably, objective task performance showed no significant association with either full trust measure (suspect accuracy: r = .18, -.20, p > .35; overall accuracy: r = -.05 to .19, p > .37), but did show a significant negative association with the TI-HRC Reliability subscale (r = .40, p < .05), as well as a negative association with detected frustration, silent periods and communication breakdowns. The dissociation between task outcomes and overall trust ratings suggests that participants’ trust judgments primarily reflected the quality and affective tone of the interaction process rather than instrumental task success. Task performance was not included as a covariate in primary models to avoid conditioning on a potential mediator of interaction policy effects.
@@ 4 Discussion
This pilot study examined how a responsive interaction policy influences trust during embodied, fully autonomous, spoken-language human–robot collaboration. Unlike much prior work in HRI trust research, all dialogue management, affect inference, task progression, and response generation were executed autonomously and in real time, without Wizard-of-Oz control or scripted repair. Participants were therefore exposed not only to adaptive interaction behaviour, but also to the latency, recognition errors, and coordination failures characteristic of deployed autonomous systems. The study was conducted as a pilot, with the primary aim of assessing feasibility and clarifying mechanisms under these realistic constraints rather than providing definitive confirmatory tests.
@@
Trust in human–robot interaction is widely recognized as multidimensional, encompassing perceptions of reliability and competence as well as affective comfort and perceived collaboration [24, 13]. The use of two validated trust measures in the present study allowed these components to be examined in parallel. Rather than treating trust as a static attribute of the system, the results suggest that it functioned as an evaluation of the collaborative experience as it unfolded over time. Notably, affective and experiential components of trust were more sensitive to interaction degradation than reliability-focused evaluations, reinforcing the view that trust in autonomous HRI is shaped by process-level interaction dynamics rather than performance outcomes alone.
@@ 5 Limitations
Measurement-related factors may have introduced additional noise. One trust instrument relied on continuous slider-based responses administered via a laptop touchpad, whereas the other used discrete, clickable Likert-style responses. Touchpad-based slider interaction can be awkward and imprecise for some users, which may have attenuated effects on the continuous scale relative to the Likert-based measure. Future work should consider standardizing response modalities across trust measures to reduce measurement noise and improve sensitivity.
@@ 6 Conclusions and Future Work
This pilot study provides evidence that affect-responsive interaction policies can influence trust during fully autonomous, in-person human–robot collaboration. The observed effects emerged independently of overall task accuracy, reinforcing the central role of interaction quality and affective responsiveness in shaping collaborative experience.

### v2

@@ 2.6 Analytic Strategy
Analyses were structured to explicitly account for interaction viability, given the spoken-language demands of the task and the variability in communication quality observed across sessions (see Section 2.5). Sessions were classified as non-viable when severe communication breakdown prevented sustained dialogue between the participant and the robot, rendering the experimental manipulation inoperative.
Communication viability was operationalized using a dialogue-level metric derived from manual coding of system logs. For each session, the proportion of dialogue turns affected by speech-recognition failure (i.e., fragmented or unintelligible utterances) was computed. For the purposes of dialogue coding, a turn was defined as a participant utterance and the robot’s subsequent response, treated as a single exchange unit. Sessions in which more than 60% of dialogue turns were affected were classified as non-viable, reflecting cases in which spoken-language interaction could not be meaningfully sustained. This criterion closely aligned with sessions independently flagged by the researcher during administration. It should be noted that because communication viability may itself be influenced by interaction policy, excluding non-viable sessions may have introduced selection bias and inflated estimated policy effects.
For this reason, analyses were conducted using three complementary approaches. Primary analyses were performed on an eligible sample excluding non-viable sessions, reflecting interactions in which the spoken-language protocol and experimental manipulation operated as intended. Full-sample analyses including all sessions were conducted as sensitivity checks. Finally, mechanism-focused analyses compared viable and non-viable sessions on interaction-process metrics (e.g., ASR failure rates, dialogue turn completion, task abandonment) to characterise how severe communication breakdown alters interaction dynamics. Trust measures from non-viable sessions were not interpreted as valid estimates of human-robot trust under functional interaction, as the robot was unable to sustain dialogue or collaborative behaviour in these cases.
All analyses were conducted using R (version 4.5.1) within the Quarto framework. Data manipulation and visualization utilized the tidyverse suite of packages [40], with mixed-effects models fitted using the lme4 and lmerTest packages [3, 23] while Bayesian hierarchical models were fitted using the brms package [6]. Summary tables were generated using the gtsummary package [36].
@@ 3 Results
Of the 29 completed sessions, 5 met the pre-specified criterion for non-viable interaction due to severe and persistent communication failure. These sessions were characterized by high rates of ASR failure, incomplete dialogue sequences, and skipped task stages. Primary results are therefore reported for the eligible sample (n = 24), with full-sample and mechanism-focused analyses reported as sensitivity and exploratory analyses, respectively.
3.1 Descriptive Analysis: Eligible Sample
Simple descriptive comparisons of post-interaction trust measures indicated higher trust ratings in the responsive condition relative to the control condition across both trust scales (see Table 2). As Figure 3 shows, average raw post-interaction scores on the Trust in Industrial Human-Robot Collaboration scale (TI-HRC) differed by approximately 26 points (Likert 1-5 converted to 0-100 scale for easier comparison across scales), while differences on the Trust Perception Scale-HRI (TPS-HRI) were approximately 15 points higher in the responsive condition compared to the control. Overall task accuracy did not differ significantly between conditions (detailed task performance results are reported below), suggesting that observed differences in trust were not driven by differential task success but rather by the quality of the interaction process itself.
Table 2: Post-Interaction Trust and Task Accuracy (eligible sample)
@@ 3.1 Descriptive Analysis: Eligible Sample
Analysis of dialogue interaction patterns confirmed successful manipulation of the interaction policies (see Table 3). Manual coding of dialogue transcripts revealed substantial and significant differences in robot behaviour across conditions (see Appendix C for coding scheme). In the responsive condition, the robot employed encouragement in 36% of dialogue turns (e.g., “You’re asking great questions!”) compared to 0% in the control condition (p < .001), expressed empathy or acknowledged participant affect in 13% of turns (e.g., “I sense you’re frustrated, but we can figure this out together!”) versus 0% in the control condition (p < .001), and used collaborative language (e.g., “we,” “let’s”) in 42% of turns compared to 5% in the control group (p < .001). While both conditions included proactive check-ins following periods of participant silence, the control robot engaged in such check-ins at a higher rate (21% vs. 13% of turns, p = .033), which reflects both the lower levels of collaborative engagement from participants resulting in more silence and the control condition’s reactive policy that limited assistance to structured silence monitoring.
@@
Importantly, the proportion of turns coded with communication breakdowns (i.e., sentence fragments or complete misunderstandings) did not differ between conditions (25% vs. 22%, p = .70), indicating speech recognition viability among eligible participants was the same between groups. Participants in the responsive condition also exhibited higher levels of AI-detected engagement during interaction (i.e., the AI “perceived” participants as responsive and interactive in collaboration), with an average of 3.50 engaged responses (SD = 1.95) compared to 2.00 (SD = 2.21) in the control condition. Consistent with these dialogue differences, interactions in the responsive condition were characterized by longer session durations and slower robot response times which reflects the latency and time associated with executing additional dialogue and affective support behaviours. As such, responsiveness in the present study reflects a bundled interaction profile rather than a single isolated manipulation. Observed trust differences may therefore partially reflect differences in interaction exposure or dialogue richness in addition to affect-adaptive behaviour. Future work should more carefully disentangle these components.
@@ 3.2 Hierarchical Models
Bayesian models used weakly informative, scale-appropriate priors. The intercept was assigned a Normal(50, 25) prior, reflecting the 0-100 trust score range and centering estimates near the midpoint of the scale. Fixed effects were assigned Normal(0, 10) priors to constrain effects to plausible magnitudes. Random-effect and residual standard deviations were assigned Exponential(1) priors to regularize variance estimates while retaining flexibility. All models demonstrated satisfactory convergence ($\hat{R} \leq$ 1.01; effective sample sizes > 1000).
@@ 3.2.1 Primary Analysis: Eligible Sample (n = 24)
Bayesian estimation converged on similar effect magnitudes with high posterior certainty. As illustrated in Figure 4, for TI-HRC, the responsive condition showed a posterior median effect of $\beta$ = 14.98 (95% credible interval [7.29, 22.22]), with near-unity probability of a positive effect. For TPS-HRI, the posterior median was $\beta$ = 12.76 (95% credible interval [2.96, 22.06]), with posterior probability exceeding 99% that the effect was positive. As expected baseline NARS showed credible negative associations with both outcomes, while native English fluency showed negative associations that were credible for TPS-HRI but uncertain for TI-HRC. Model fit was substantial for TPS-HRI (conditional $R^{2}$ = 0.64) and moderate for TI-HRC (conditional $R^{2}$ = 0.42), with fixed effects explaining 16% and 21% of variance, respectively. The smaller item-level variance for TI-HRC suggests greater coherence among affective trust items under functional interaction conditions.
Figure 4: Posterior distributions of fixed effects from the final Bayesian mixed-effects models (eligible sample: n=24) predicting post-interaction trust. Fixed effects include policy, Negative Attitudes Towards Robots (NARS) and language (binary). Half-eye densities show posterior distributions; points indicate posterior medians; thick and thin intervals denote 80% and 95% credible intervals, respectively. The dashed vertical line marks a null effect, and positive values indicate higher trust scores on the respective scale (points on a 0-100 scale).
@@ 3.2.3 Mechanism Analysis: Communication Breakdown (n=29)
To examine whether communication quality moderated the effect of interaction policy, Bayesian models were fitted in the full sample with the proportion of turns characterized by communication breakdown included as an interaction term. Predicted trust as a function of breakdown proportion for each condition and trust measure is shown in Figure 5.
Figure 5: Predicted trust as a function of communication breakdown by interaction policy. Panels show experienced trust (left) and perceived trust (right). Lines represent posterior medians and shaded bands indicate 80% credible intervals. Communication breakdown is defined as the proportion of interaction turns flagged as breakdowns; predictions are shown at mean baseline negative attitudes toward robots.
The two trust dimensions responded differently to communication degradation. For experienced trust (TI-HRC), the responsive condition showed consistently higher trust than control at low breakdown levels, but this advantage attenuated as breakdown increased (posterior median $\beta$ = -5.97, 95% CI [-23.01, 10.92]; 76% probability negative). This pattern is consistent with the interpretation that experienced trust depends on the robot’s sustained ability to deliver responsive behaviour — when communication fails frequently, the affect-adaptive features that define the responsive policy cannot be expressed, and the trust advantage erodes. In contrast, perceived trust (TPS-HRI) showed a weak, unstable interaction centered near zero (posterior median $\beta$ = -1.14, 95% CI [-18.87, 16.28]; 55% probability negative), suggesting evaluative reliability judgements are less sensitive to graded communication degradation; at least in this small sample. This may reflect the fact that perceived trust is anchored by discrete moments of successful collaboration: a robot that continues attempting to engage despite communication difficulties may still signal competence, even if sustained rapport cannot be established.
These results clarify the sensitivity analysis findings. The attenuation of the responsive policy effect in the full sample reflects the inclusion of sessions in which the communication level required to deliver the responsive manipulation had substantially failed (defined as >60% of turns affected by ASR failure) and is not simply noise from increased variability. This supports both the exclusion criterion applied in the primary analysis and the broader interpretation that experienced trust, in particular, is contingent on sustained interaction quality. Put plainly, a robot that cannot understand you cannot adapt to you and without adaptation, the responsive policy offers nothing the control policy does not.
@@ 3.4 Individual differences and correlational patterns
Consistent with the experimental manipulation, specific robot dialogue behaviours showed substantial correlations with trust outcomes. Robot empathy expression—quantified as the proportion of dialogue turns in which the robot acknowledged participant affect or expressed understanding (e.g., “I sense your frustrated, how can I help?”)—was strongly correlated with both experienced trust (TI-HRC; r = .53, p = .008) and perceived trust (TPS-HRI: r = .65, p = .001). Similarly, robot use of collaborative language (e.g., “we,” “let’s”) was positively associated with experienced trust (TI-HRC: r = .50, p = .012), as was robot encouragement (TI-HRC: r = .45, p = .027; TPS-HRI: r = .42, p = .040). These associations suggest that the specific affective and collaborative behaviours implemented in the responsive condition were linked to participants’ trust evaluations beyond simple condition assignment.
@@
Notably, objective task performance showed no significant association with either full trust measure (suspect accuracy: r = .18-.20, p > .35; overall accuracy: r = -.05 to .19, p > .37), but did show a significant negative association with the TI-HRC Reliability subscale (r = .40, p < .05), as well as a negative association with detected frustration, silent periods and communication breakdowns. The dissociation between task outcomes and overall trust ratings suggests that participants’ trust judgments primarily reflected the quality and affective tone of the interaction process rather than instrumental task success. Task performance was not included as a covariate in primary models to avoid conditioning on a potential mediator of interaction policy effects.
@@ 4 Discussion
This pilot study examined how a responsive interaction policy influences trust during embodied, fully autonomous, spoken-language human-robot collaboration. Unlike much prior work in HRI trust research, all dialogue management, affect inference, task progression, and response generation were executed autonomously and in real time, without Wizard-of-Oz control or scripted repair. Participants were therefore exposed not only to adaptive interaction behaviour, but also to the latency, recognition errors, and coordination failures characteristic of deployed autonomous systems. The study was conducted as a pilot, with the primary aim of assessing feasibility and clarifying mechanisms under these realistic constraints rather than providing definitive confirmatory tests.
@@
Trust in human-robot interaction is widely recognized as multidimensional, encompassing perceptions of reliability and competence as well as affective comfort and perceived collaboration [24, 13]. The use of two validated trust measures in the present study allowed these components to be examined in parallel. Rather than treating trust as a static attribute of the system, the results suggest that it functioned as an evaluation of the collaborative experience as it unfolded over time. Notably, affective and experiential components of trust were more sensitive to interaction degradation than reliability-focused evaluations, reinforcing the view that trust in autonomous HRI is shaped by process-level interaction dynamics rather than performance outcomes alone.
@@ 5 Limitations
Measurement-related factors may have introduced additional noise. The TPS-HRI trust instrument (perceived trust) relied on a continuous slider-based input administered via a laptop touchpad, whereas the TI-HRC relied on a discrete, clickable, Likert-style response. Touchpad-based slider interaction can be awkward and imprecise for some users, which may have attenuated effects on the continuous scale relative to the Likert-based measure. Future work should consider standardizing response modalities across trust measures to reduce measurement noise and improve sensitivity as well as incorporating additional process-level measures of trust dynamics during interaction (e.g., real-time trust ratings, physiological indicators) to complement post-interaction evaluations.
@@ 6 Conclusions and Future Work
This pilot study provides evidence that affect-responsive interaction policies can influence trust during fully autonomous, in-person human-robot collaboration. The observed effects emerged independently of overall task accuracy, reinforcing the central role of interaction quality and affective responsiveness in shaping collaborative experience.
@@
7 Ethical Approval
All procedures were approved by the Laurentian University Research Ethics Board (ROMEO File: #6021966). Participants provided written informed consent prior to participation and were debriefed in-person following the session. Data were de-identified and stored securely in accordance with institutional guidelines.
8 Data and Code Availability
The analysis code (R) is available at: GitHub Repository. The interaction system and prompt templates used to implement the experimental conditions (Python) are also available in the same repository. De-identified participant data (survey scales and session-level summary measures) are available upon reasonable request to the corresponding author. Raw interaction logs are not publicly available due to participant privacy constraints.
9 Conflict of Interest
The authors declare no conflicts of interest.
10 Author Contributions
S. Heron: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization.
M.C. Lau: Conceptualization, Methodology, Writing - Review & Editing, Supervision.
11 Funding
This work was supported by IAMGOLD President’s Innovation Fund (Grant No. 2025-20377).
12 Declaration of Generative AI and AI-assisted Technologies in the Writing Process
During the preparation of this work, the author(s) used Claude (Sonnet 4.6, Anthropic) to review the manuscript for clarity and readability. After reviewing the AI-generated suggestions, the author(s) made revisions to the text as appropriate. The author(s) takes full responsibility for the content of the published work.